OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. please help Set #1 Set #2 (0, 4) and (-2, 8) (3, 7) and (-1, -5)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

do you know how to find the slopes of the points ?

OpenStudy (anonymous):

why did you close the question ? Do you still need help ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok :) bye for now

OpenStudy (anonymous):

oh yes i need help

OpenStudy (anonymous):

on this question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets find the slopes using the slope formula : slope(m) = (y2 - y1)/(x2 - x1) set 1 (0,4) x1 = 0 and y1 = 4 (-2,8) x2 = -2 and y2 = 8 lets sub our info in : m = (8 - 4) / (-2 - 0) m = 4/-2 m = -2 <-- slope you see how I found the slope for set 1 ?

OpenStudy (anonymous):

yes but it says The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

now we need to find the slope of set 2 and compare them. m = (y2 - y1) / (x2 - x1) (3,7) x1 = 3 and y1 = 7 (-1,-5) x2 = -1 and y2 = -5 m = (-5 - 7) / (-1 - 3) m = -12/-4 m = 3 slope for set 1 is -2 and slope for set 2 is 3 so what do you think the answer is ?

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

you are correct :)

OpenStudy (anonymous):

good job :)

OpenStudy (anonymous):

do you need anymore help ?

OpenStudy (anonymous):

thank u and yes i do

OpenStudy (anonymous):

what do you need help with ?

OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation x + 3y = 3 using the x- and y- intercepts method? Plot the y-intercept (0, 1). Plot the intercept (3, 0). Count down 1 and right 3 from the y-intercept. Draw a line through the two points

OpenStudy (anonymous):

i think its b

OpenStudy (anonymous):

i mean c

OpenStudy (anonymous):

it is C

OpenStudy (anonymous):

i got What is the slope of the line 2x - 5y = -10? negative five halves five halves two fifths negative two fifths none of the above

OpenStudy (anonymous):

nd i think its none of the above

OpenStudy (anonymous):

2x - 5y = 10 -5y = -2x + 10 (multiply by -1 to make y positive) 5y = 2x - 10 (now divide by 5) y = 2/5x - 2 y = mx + b and m is your slope, therefore your slope is 2/5

OpenStudy (anonymous):

wait...I wrote the problem down wrong...hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

2x - 5y = -10 -5y = -2x - 10 5y = 2x + 10 y = 2/5x + 2 it didn't matter, the slope is still 2/5

OpenStudy (anonymous):

Explain how to graph 2x + y = 14 using the x- and y-intercepts method. Be sure to use complete sentences and use the x- and y-intercepts method only.

OpenStudy (anonymous):

to find the x intercept, sub in 0 for y. 2x + y = 14 2x + 0 = 14 2x = 14 x = 7 <-- x intercept = (7,0) to find the y intercept, sub in 0 for x 2x + y = 14 2(0) + y = 14 y = 14 <-- y intercept = (14,0) so you have a point on the x axis at 7 and a point on the y axis at 14. Now just draw a line connecting the two points.

OpenStudy (anonymous):

thank yu so much

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

;] your the best

OpenStudy (anonymous):

I don't know about that...but thank u anyways :)

OpenStudy (anonymous):

lol its tru